# Layered Multiple Scattering Approach to Hard X-ray Photoelectron Diffraction: Theory and Application

### v1

@@ Abstract
Photoelectron diffraction (PED) is a powerful and essential experimental technique for resolving the structure of surfaces with sub-angstrom resolution. In the high energy regime, researchers in angle-resolved photoemission spectroscopy (ARPES) observe modulating patterns attributed to X-ray-PED (XPD) effects. This is accompanied by other challenges such as low cross-sections, significant photon momentum transfer, and non-negligible phonon scattering. Overall, XPD is not only an advantageous approach but also exhibits unexpected effects. To disentangle these diffraction influences, we present a PED implementation for the SPRKKR package that utilizes multiple scattering theory and a one-step model in the photoemission process. Unlike real-space implementations of the multiple scattering XPD formalism, we propose a k-space implementation based on the layer KKR method. The main advantage of this method is its ability to address a very broad kinetic energy range (20-8000 eV) without convergence problems related to angular momentum and cluster size. Furthermore, the so-called alloy analogy model can be used to simulate XPD at finite temperatures as well as XPD effects observed in soft and hard X-ray ARPES. For practical applications, we have calculated the circular dichroism in angular distributions (CDAD) associated with core-level photoemission of 2p from Si(100) and 3p from Ge(100). Photoelectrons are excited by hard X-rays (6000 eV) with right and left circularly polarized radiation (RCP and LCP, respectively).
@@ 1 Introduction
As an influential technique for probing the atomic structure in the vicinity of a given emitter, photoelectron diffraction (PED) is widely used for various investigation purposes: crystal structures, bonding geometries of atoms, and the local environment of impurity or dopant atoms inside surfaces [1, 2, 3, 4]. It is regarded as being analogous to angle-resolved photoemission spectroscopy (ARPES) at the fundamental level, specifically regarding the angular distribution of photoelectrons emitted from a crystal surface. Nonetheless, the underlying physics and the investigative objectives of the two approaches are disparate. The angular distribution of emitted electrons represents the momentum of initial states in ARPES while in PED, it reveals the interference of photoelectron waves from final states. The observed alteration in PED intensities depends on the constructive or destructive interference of all the coherently emitted electrons. This means that details related to atomic arrangements, bonding orientations and distances, and electronic and chemical features are analyzed. Depending on the utilized photon energies with respect to photoelectron kinetic energies, this tool can be termed either ultraviolet-PED (UPD) or X-ray-PED (XPD). In summary, photoemission can provide information about core-level electrons or delocalized valence band electrons. Accordingly, there are core-level PED (CL-PED) and valence-band PED (VB-PED). The main difference is that the angular distribution in VB-PED is induced by two interference phenomena: from primary waves emitted at various atomic sites and from the scattering of corresponding secondary waves at neighboring sites [5]. At high photon energies, XPD effects [6, 7, 8, 9, 10, 11, 12] make ARPES data vague, alongside other complications such as poor cross-sections, considerable momentum transfer from photons causing a pronounced modulation of ARPES patterns, and substantial phonon scattering.
@@
Given the significant contributions of PED (especially in the circular dichroism in the angular distribution (CDAD) interpretation and disentangling of XPD effects), the development of a suitable theoretical model of scattering is of paramount importance. Powerful computational approaches that address single and multiple scattering in real space and reciprocal space have been developed earlier and will be summarized and reviewed in subsection 2.1. However, these packages share some limitations. First, most of them perform calculations in real space [13, 14, 15], which might not be convenient for high kinetic energies where the effective cluster sizes need to be very large. Second, some models have to borrow the atomic potential from other sources. For instance, in the MSCD code [16], it is taken from the muffin-tin potential in the database of Moruzzi et al. [17], which neglects relativistic effects. For low kinetic energy, MsSpec [18] utilizes the potential file from Munich SPRKKR [19, 20]. In the case of the dynamical high-energy electron diffraction approach to XPD [21, 22], the real and imaginary parts of the crystal potential (assumed to have 3D bulk periodicity) are computed based on the Fourier coefficients of atomic potentials available in various parameterizations [23]. Third, some approaches are designed to cope with a specific regime of kinetic energy, not flexibly covering the whole range from UV to hard X-ray. For instance, EDAC [24] is designed to cope with diffraction at lower kinetic energies compared to the method proposed by Winkelmann et al. [21, 22]. Fourth, although hard X-ray calculations (e.g. a chromium Kα source) [25, 26] attempt to simulate XPD patterns, not all of them are available for CDAD patterns for core levels in this regime, to the best of our knowledge. Indeed, the TMSP code [15] developed by Matsushita and colleagues can work with circularly polarized light [27] and interpret the CDAD in terms of the rotational shift of forward focusing peaks around the incident-light axis [28, 29]. However, no simulated CDAD for hard X-ray core levels has been published. Fifth, Kikuchi diffraction [30] is well-documented in scanning and transmission electron microscopy (SEM and TEM) with energy sources up to 100 keV, where it manifests as electron backscatter diffraction patterns [31]. Nevertheless, it is challenging to observe this phenomenon in spectroscopy and few simulations are able to reconstruct it [32, 33, 34]. Lastly, to accurately represent scattering by cluster approaches, there is a high need for large values for the maximum angular momentum $l_{max}$ (the number of scattering phase shifts). In the case of 10 keV electrons scattering from an atomic potential with a radius of 1 Å radius, this parameter must be raised to 100 [22].

### v3

@@ Abstract
Photoelectron diffraction (PED) is a powerful technique for resolving surface structures with sub-angstrom precision. At high photon energies, angle-resolved photoemission spectroscopy (ARPES) reveals PED effects, often challenged by small cross-sections, momentum transfer, and phonon scattering. X-ray PED (XPD) is not only an advantageous approach but also exhibits unexpected effects. We present a PED implementation for the spin-polarized relativistic Korringa-Kohn-Rostoker (SPRKKR) package to disentangle them, employing multiple scattering theory and a one-step photoemission model. Unlike conventional real-space approaches, our method uses a k-space formulation via the layer-KKR method, offering efficient and accurate calculations across a wide energy range (20–8000 eV) without angular momentum or cluster size convergence issues. Additionally, the alloy analogy model enables simulations of finite-temperature XPD and effects in soft/hard X-ray ARPES. Applications include modeling circular dichroism in angular distributions (CDAD) in core-level photoemission of Si(100) 2p and Ge(100) 3p, excited by 6000 eV photons with circular polarization.
@@ 1 Introduction
As an influential technique for probing the atomic structure in the vicinity of a given emitter, photoelectron diffraction (PED) is widely used for various investigation purposes: crystal structures, bonding geometries of atoms, and the local environment of impurity or dopant atoms inside surfaces [1, 2, 3, 4]. It is considered analogous to angle-resolved photoemission spectroscopy (ARPES) at the fundamental level, specifically regarding the angular distribution of photoelectrons emitted from a crystal surface. However, the underlying physics and the investigative objectives of the two approaches are different. The angular distribution of the emitted electrons represents the momentum of the initial states in ARPES, while in PED it reveals the interference of photoelectron waves from final states. The observed change in PED intensities depends on the constructive or destructive interference of all the coherently emitted electrons. This means that details related to atomic arrangements, bonding orientations and distances, and electronic and chemical properties are analyzed. Depending on the utilized photon energies with respect to photoelectron kinetic energies, this tool can be termed either ultraviolet-PED (UPD) or X-ray-PED (XPD). In summary, photoemission can provide information about core-level electrons or delocalized valence band electrons. Accordingly, there are core-level PED (CL-PED) and valence-band PED (VB-PED). The main difference is that the angular distribution in VB-PED is caused by two interference phenomena: from primary waves emitted at different atomic sites and from the scattering of corresponding secondary waves at neighboring sites [5]. At high photon energies, XPD effects [6, 7, 8, 9, 10, 11, 12] obscure ARPES data, along with other complications such as poor cross-sections, considerable momentum transfer from photons causing a pronounced modulation of ARPES patterns, and substantial phonon scattering.
@@
Given the significant contributions of PED (especially in the circular dichroism in the angular distribution (CDAD) interpretation and disentangling of XPD effects), the development of a suitable theoretical model of scattering is of paramount importance. Powerful computational approaches that address single and multiple scattering in real space and reciprocal space have been developed earlier and will be summarized and reviewed in the section 4.1. However, these packages share some limitations. First, most of them perform calculations in real space [13, 14, 15], which might not be convenient for high kinetic energies where the effective cluster sizes need to be very large. Second, some models have to borrow the atomic potential from other sources. For instance, in the MSCD code [16], it is taken from the muffin-tin potential in the database of Moruzzi et al. [17], which neglects relativistic effects. For low kinetic energy, MsSpec [18] utilizes the potential file from Munich SPRKKR [19, 20]. In the case of the dynamical high-energy electron diffraction approach to XPD [21, 22], the real and imaginary parts of the crystal potential (assumed to have 3D bulk periodicity) are computed based on the Fourier coefficients of atomic potentials available in various parameterizations [23]. Third, some approaches are designed to cope with a specific regime of kinetic energy, not flexibly covering the whole range from UV to hard X-ray. For instance, EDAC [24] is designed to describe diffraction at lower kinetic energies in comparison to the method proposed by Winkelmann et al. [21, 22]. Fourth, although hard X-ray calculations (e.g. a chromium Kα source) [25, 26] attempt to simulate XPD patterns, not all of them are available for CDAD patterns for core levels in this regime, to the best of our knowledge. Indeed, the TMSP code [15] developed by Matsushita and colleagues can work with circularly polarized light [27] and interpret the CDAD in terms of the rotational shift of forward focusing peaks around the incident-light axis [28, 29]. However, no simulated CDAD for hard X-ray core levels has been published. Fifth, Kikuchi diffraction [30] is well-documented in scanning and transmission electron microscopy (SEM and TEM) with energy sources up to 100 keV, where it manifests as electron backscatter diffraction patterns [31]. Nevertheless, it is challenging to observe this phenomenon in photoemission spectroscopy, and few simulations are able to reconstruct it [32, 33, 34]. Lastly, to accurately represent scattering by cluster approaches, there is a high need for large values for the maximum angular momentum $l_{max}$ (the number of scattering phase shifts). In the case of 10 keV electrons scattering from an atomic potential with a radius of 1 Å radius, this parameter must be increased to 100 [22].